OpenStudy (anonymous):

Solve the system of equations: what is the y-value in the solution? 3x - 5y = -16 and 2x + 5y = 31

OpenStudy (anonymous):

\(3x - 5y = -16\)..........................1 multiplying by 3 \(2x + 5y = 31\)............................2multiplying by -2 6x-15y=-48................1 -6x-10y=-62...............2 add equation 1 to 2 -25 y =-110 y = 4.4 if you want to find x .. 2x-5y=-16 2x-(5× 4.4)=-16 2x=6 x= 3 You are welcome

OpenStudy (anonymous):

Thanx for replying. yuppy i got the correct answer.you're a life saver!

OpenStudy (anonymous):

um thats not one of the multiple question answers.

OpenStudy (anonymous):

so your answer is 31/5? correct?

OpenStudy (anonymous):

lool. no worries

OpenStudy (anonymous):

its a fraction

OpenStudy (anonymous):

aha, i see but i thought you meant to say 5 not 31/5

OpenStudy (anonymous):

your method is implement in first equation not on second one .. and i very confused about how you choosed x = 5 ? any way .. 2 mints solution = no wasting time .. thnx for sharing your ideas .

OpenStudy (anonymous):

ya i know sorries

OpenStudy (anonymous):

it's ok. the answer was 3. but i see you are trying to make effort :D we all make mistakes in math.